Consider the fission reaction 1



0n + 235



92U →



89



37Rb + 3 0



-1e + 3 1



0n + 144



58Ce



How much energy in kilojoules is released per gram of Uranium- 235?

1
Expert's answer
2022-02-08T09:13:02-0500

To determine the E? value of a cell, we determine the anode and cathode half cell reactions and then calculate the E? by the help of the formula:


E ? ( cell ) = E? (cathode) - E? (anode)


The half-cell reactions are written as below:


Anode: Na(s) → Na+ (aq) + e- E? (anode) = -2.714V


Cathode: Fe2+ (aq) + 2e- → Fe(s) E? (cathode) = -0.409V


So, E? (cell) = -0.409 – (-2.714) = 2.305V